OpenStudy (anonymous):

Evaluate the integral ∫0 (top) sqrt(t+1)dt 7 (bottom) ok

OpenStudy (anonymous):

I on't get it. Its soo confusing the way you have it.

OpenStudy (anonymous):

I will retype it again

OpenStudy (anonymous):

\[\int_7^0\sqrt{t+1}dt\] maybe?

OpenStudy (anonymous):

7 is on top 0 is on the bottom. Sorry it was backwards earlier.

OpenStudy (anonymous):

You must be in 10 or 11th grade. I am only in 9th lol.

OpenStudy (anonymous):

\[\int_0^7\sqrt{t+1}dt\] anti derivative is \[\frac{2}{3}(x+1)^{\frac{3}{2}}\] by the power rule backwards evaluate at \(7\) and at \(0\) then subtract

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw